The product -3^4 * 3^3 =
(1) 3^7
(2) -3^7
(3)9^7
(4)-9^7

Answer by bucky(2189)

You are asked to find the product:
.
-3^4 * 3^3
.
Note -3^4 does not mean (-3)^4 which would be -3*-3*-3*-3
.
Instead -3^4 means -(3*3*3*3) = -(3^4)
.
Therefore you can think of the given product as:
.
-1*(3^4)*(3^3)
.
in this case, the two terms in parentheses both have the base 3. Therefore, they can be multiplied
by adding their exponents ... making the product:
.
-1*(3^(4+3)) = -1*3^7 = -3^7
.
So the correct answer to this problem is choice b) in your list of answers.
.
Hope this helps you to understand the problem.
.